The age of Rida will be 6 times after 24 years. What is the present age of Rida?
Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

let x equal how old rida is now.
in 24 years rida will be x + 24 years old.
that's 6 times as old as rida is now.

the formula for that is x + 24 = 6x
subtract x from both side of that equation to get 24 = 5x.
solve for x to get:
x = 24/5 = 4.8

rida is 4.8 years old now.
rida will be 4.8 + 24 years old in 24 years = 28.8
28.8 / 4.8 = 6 means that, in 24 years, rida will be 6 times as old as rida is now.
this confirms the solution is correct.

Answer by MathTherapy(10552)


The age of Rida will be 6 times after 24 years. What is the present age of Rida?
Doesn't make sense since her calculated age is NOT an INTEGER!